OpenStudy (anonymous):

Find all integer zeros of the function. x^4 + x^3 - 3x^2 - x + 2 = 0

OpenStudy (amoodarya):

x^4 + x^3 - 3x^2 - x + 2 =x^4 + x^3 - 2x^2 - x-x^2 + 2 =x^3(x+1) -2(x^2-1) -x(1+x) =(x+1)(x^3-2(x-1)-x) can you go on ?

OpenStudy (amoodarya):

=(x+1)(x^3-3x+2) =(x+1)(x^3-x-2x+2) =(x+1)(x(x^2-1)-2(x-1)) =(x+1)(x-1)(x(x+1)-2)

OpenStudy (anonymous):

the roots would be -1, +1, +1 and -2?

OpenStudy (amoodarya):

yes

OpenStudy (anonymous):

thank you so much!